Find the rate of change between the two points: (8, 1) and (4, 2) where x is in months and y is in inches.
Answer by tommyt3rd(5050)

Rate of change is computed using slope:
m=(y2-y1)/(x2-x1)
so
m=(2-1)/(4-8)=-1/4
that means the rate is falling at 0.25 inches every month